OpenStudy (anonymous):

Find the altitude of an equilateral triangle, if each side of the triangle has a length of 14 meters

OpenStudy (kkutie7):



OpenStudy (kkutie7):

a^2+7^2=14^2 a^2=14^2-7^2 a^2=147 a= square root of 147 a=about 12.12